OpenStudy (anonymous):

Solve 3x2 + 4x = 2. quantity of 2 plus or minus square root of 10 all over 6 quantity of negative 2 plus or minus 2 square root of 10 all over 3 quantity of negative 2 plus or minus square root of 10 all over 3 quantity of 4 plus or minus square root of 10 all over 3

OpenStudy (solomonzelman):

\[3x^2 + 4x = 2\]\[3x^2+4x-2=0\]use the quadratic formula.

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

im still confused

OpenStudy (solomonzelman):

\[\frac{-b±\sqrt{b^2-4ac}}{2a}\] a=Coefficient next to x^2 b= Coefficient next to x c=Constant Always set the equation equal to zero before applying the above formula- (the quadratic formula) just like I did.